OpenStudy (anonymous):

Solve by completing the square 2x^2 -4x -4=0 A. 1+- √3 B.-1+- √3 C.-1+- 2√3 D.1 +- 2√3

OpenStudy (campbell_st):

divide every term by 2 and you get \[x^2 -2x - 2 = 0\] rewrite the equation as \[x^2 -2x + ?= 2 +?\] what do you need to add to both sides of the equation to get a perfect square on the left

OpenStudy (wolf1728):

2x^2 -4x -4=0 To complete the square: Move the "non X" term to the right: 2x^2 -4x = 4 Divide the equation by the coefficient of X² x² -2x = 2 take the coefficient of X which is -2 divide that by 2 which is -1 square that number +1 then add it to both sides of the equation x² - 2x +1 = 3 Take the square root of BOTH sides of the equation x-1 = sqrt(3) x = sqrt(3) +1 x = 1.732 +1 x = 2.732 sqrt(3) also = -1.732 x = -1.732 +1 x = -.732